OpenStudy (anonymous):

Find the open intervals on which the function is concave up or concave down f(x)=1/ x^2 + 3 WILL GIVE MEDAL PLEASE HELP!

OpenStudy (anonymous):



OpenStudy (anonymous):

what did you get for the first derivative?

OpenStudy (anonymous):

X^2+3-2x/(x^2+3)^2

OpenStudy (anonymous):

Where did you go @satellite73 ?

zepdrix (zepdrix):

Hmm I'm a little worried about your first derivative. Let's simply use the power rule to do this one.\[\large f(x)=\frac{1}{x^2+3} \qquad=\qquad (x^2+3)^{-1}\] Taking the derivative gives us,\[\large f'(x)=-(x^2+3)^{-2}(2x) \qquad=\qquad \frac{-2x}{(x^2+3)^2}\]

zepdrix (zepdrix):

Did you apply the quotient rule to do this one?

OpenStudy (anonymous):

Yeah so I guess that's why it was off

zepdrix (zepdrix):

Quotient Rule will give you the same result. Lemme show you where the mistake was.

zepdrix (zepdrix):

\[\large f'(x)=\frac{\color{royalblue}{(1)'}(x^2+3)-(1)\color{royalblue}{(x^2+3)'}}{(x^2+3)^2}\]

zepdrix (zepdrix):

So there is our setup for the quotient rule.

zepdrix (zepdrix):

What's the derivative of 1? Derivative of a constant, hmmm.

OpenStudy (anonymous):

LOL oops! It's been a long day and I'm just so stressed :/

zepdrix (zepdrix):

:3

OpenStudy (anonymous):

So now what do I do once I find the derivative?

zepdrix (zepdrix):

\[\large f'(x)=0 \qquad\to\qquad \text{critical points}\]When we look at test points around the critical points, it tells us where the function is `increasing and decreasing`. \[\large f''(x)=0 \qquad\to\qquad \text{inflection points}\]When we look at test points around the inflection points, it tells us where the function is `concave up and concave down`. So we'll want to take the derivative once more, then set it equal to zero.

zepdrix (zepdrix):

So looks like we'll need to apply the Quotient Rule again.

zepdrix (zepdrix):

So here is our setup, \[\large f''(x)=\frac{\color{royalblue}{(-2x)'}(x^2+3)^2-(-2x)\color{royalblue}{\left[(x^2+3)^2\right]'}}{\left[(x^2+3)^2\right]^2}\]